OpenStudy (lexy148):

Olivia needs to determine the approximate height of a tree in her yard. she measures the length of the shadow of the tree to be 20ft. at the same time, standing beside the tree, her shadow is 10.5ft long and her height is 5ft 4in. approximately tall is the tree?

OpenStudy (math&ing001):

 \[\frac{ 5 ft 4 inches. }{ 10.5ft }=\frac{ treeheight }{ 20ft } \]

OpenStudy (anonymous):

\[\frac{10+\frac{5}{10}}{5+\frac{4}{12}}=\frac{20}{h} \]\[h=\frac{640}{63}=10.1587 \]